OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Part 1: In complete sentences, explain the necessary steps required to find the geometric partial sum the sum (Look at equation below). Be sure to give your final answer in your explanation. (3 points) Part 2: Is this partial sum divergent or convergent in nature? Explain your answer.

OpenStudy (magbak):

\[\sum_{i=1}^{5}3(0.1)^(i-2)\]

OpenStudy (magbak):

@CarlosGP @primeralph @Loser66

OpenStudy (anonymous):

Now the formula is a bt diferent. If "r" is the common ratio, you have:\[S_N=\sum_{n=1}^{N}a_n=\frac{ a_1-r·a_N }{ 1-r }\]

OpenStudy (anonymous):

in your case, r=0.2 Find a1 and aN

OpenStudy (magbak):

OK one sec please.

OpenStudy (magbak):

How do I find the common ration if I am not given any terms to add and divide to reach the common ratio.

OpenStudy (anonymous):

Sorry, the common ratio in your case is 0.1. You find it in this way: \[a_{i}=3·(0.1)^{(i-2)}\] \[a_{i+1}=3·(0.1)^{(i+1-2)}=3·(0.1)^{(i-1)}\]\[r=\frac{ a_{i+1} }{ a_i}=\frac{ 3·0.1^{i-1} }{ 3·0.1^{i-2} }=0.1\]

OpenStudy (anonymous):

got it?

OpenStudy (magbak):

OK.

OpenStudy (anonymous):

now apply the formula

OpenStudy (anonymous):

a1=? and a5=?

OpenStudy (magbak):

I do not know because I am not give a1

OpenStudy (anonymous):

But you have: ai, just replace the "i" with "1"

OpenStudy (magbak):

Oh ok

OpenStudy (anonymous):

Dont you get:\[a_1=3·0.1^{(1-2)}=3/0.1=30\]

OpenStudy (magbak):

It is 1.003

OpenStudy (anonymous):

see my calcullation and try again

OpenStudy (magbak):

ok

OpenStudy (anonymous):

now repeat for n=5

OpenStudy (anonymous):

or replace i with 5

OpenStudy (anonymous):

dont you get \[a_5=3·0.1^{5-2}=0.003\]

OpenStudy (magbak):

no

OpenStudy (anonymous):

what do you get?

OpenStudy (magbak):

Oh yes that is what I got

OpenStudy (anonymous):

excellent, now apply the formula, you have all that you need:\[S_n=\frac{30- 0.003\times 0.1 }{ 1-0.1 }=????\]

OpenStudy (magbak):

It is 33.333

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

Have you caught the whole process?

OpenStudy (magbak):

Yes but I am not sure how to sum it up how my teacher wants it written for part 1 and part 2

OpenStudy (anonymous):

We have done part1. All the process is explained above For part 2, whenever the module of the common ratio is <1 the sum is convergent

OpenStudy (magbak):

Ok. If it is not to much can you sum up all we did in one post PLESASE

OpenStudy (anonymous):

I would love to but I have to go. I think the whole process is perfectly developed throughout all the posts. This is a good exercise for you to summarize anyway!

OpenStudy (magbak):

OK can I sum it up and you tell me if it is right PLEASE

OpenStudy (anonymous):

all right but not today. i gotta go and will have a look at it tomorrow (Spain´s time). send it to me in a message

OpenStudy (magbak):

OK. Bye

OpenStudy (anonymous):

bye